# PrefixQuant: Static Quantization Beats Dynamic through Prefixed Outliers in LLMs

### v1

@@ Appendix C Details of Rotation
| (Seq_len,dimension) | Quantization Time (ms) | Speedup | |
| Per-token Dynamic | Per-tensor static | ||
| (1,4096) | 0.0127 | 0.0038 | 3.34x |
| (1,8192) | 0.0144 | 0.004 | 3.60x |
| (2048,4096) | 0.1073 | 0.0344 | 3.12x |
| (2048,8192) | 0.2084 | 0.0652 | 3.19x |
| Average Speedup | 3.31x | ||
| (Seq_len,dimension) | Quantization Time (ms) | Speedup | |
| Per-token Dynamic | Per-tensor static | ||
| (1,4096) | 0.020 | 0.0072 | 2.81x |
| (1,8192) | 0.022 | 0.0075 | 2.88x |
| (2048,4096) | 0.095 | 0.033 | 2.88x |
| (2048,8192) | 0.157 | 0.058 | 2.71x |
| Average Speedup | 2.82x | ||
| (Seq_len,input_c, output_c) | FP16 (ms) | W4A4 (ms) | ||
| Quarot | + static quant | + improved GEMV | ||
| (1,4096,4096) | 0.0512 | 0.0578 (0.89x) | 0.0472 (1.08x) | 0.0223 (2.30x) |
| (1,4096,14336) | 0.1548 | 0.0641 (2.42x) | 0.0549 (2.83x) | 0.0475 (3.27x) |
| (1,8192,8192) | 0.1080 | 0.0957 (1.77x) | 0.0863 (1.97x) | 0.0561 (3.02x) |
| (1,8192,28672) | 0.5762 | 0.2087 (2.76x) | 0.1977 (2.91x) | 0.1503 (3.83x) |
| (2048,4096,4096) | 1.0666 | 0.3699 (2.88x) | 0.2965 (3.59x) | - |
| (2048,4096,14336) | 3.5766 | 0.9358 (3.93x) | 0.8618 (4.27x) | - |
| (2048,8192,8192) | 3.9986 | 1.0211 (4.03x) | 0.8718 (4.72x) | - |
| (2048,8192,28672) | 13.1607 | 2.8609 (4.74x) | 2.7177 (4.99x) | - |
| (Seq_len,input_c, output_c) | FP16 (ms) | INT4 (ms) | ||
| Quarot | + static quant | + improved GEMV | ||
| (1,4096,4096) | 0.0418 | 0.0588 (0.71x) | 0.0455 (0.92x) | 0.0235 (1.78x) |
| (1,4096,14336) | 0.1026 | 0.0679 (1.51x) | 0.0556 (1.85x) | 0.0441 (2.33x) |
| (1,8192,8192) | 0.1080 | 0.0888 (1.22x) | 0.0735 (1.47x) | 0.0508 (2.13x) |
| (1,8192,28672) | 0.3036 | 0.1668 (1.82x) | 0.1534 (1.97x) | 0.1114 (2.72x) |
| (2048,4096,4096) | 0.2762 | 0.2408 (1.15x) | 0.1799 (1.54x) | - |
| (2048,4096,14336) | 1.0092 | 0.5461 (1.85x) | 0.4850 (2.08x) | - |
| (2048,8192,8192) | 1.0583 | 0.5298 (2.00x) | 0.4349 (2.43x) | - |
| (2048,8192,28672) | 3.6897 | 1.4686 (2.51x) | 1.3857 (2.66x) | - |
Appendix D Inference Efficiency Details
In this section, we examine the inference efficiency of PrefixQuant. We conduct tests on Nvidia RTX 3090 and A100-80GB GPUs, considering sequence lengths of 1 and 2048, with a batch size of 1. We detail the speedup ratios for quantization overhead, quantized linear layers, and end-to-end inference below.
Reduced Quantization Overhead. Activations are quantized and packed into low-bit formats for matrix manipulations. We define the time for this process during inference as quantization overhead. The per-token dynamic quantization kernel is sourced from QuaRot (Ashkboos et al., 2024b). Table 8(b) shows the speedup of per-tensor static quantization over per-token dynamic quantization. We can find that the per-tensor static quantization kernel achieves speedups of on the RTX 3090 and on the A100-80GB.
Accelerated Quantized Linear Layer. The quantized linear layer consists of quantization, low-bit matrix multiplication, and de-quantization. The speedup for the quantization process is in Table 8(b). For low-bit matrix multiplication, we use the 4-bit GEMM kernel from CUTLASS and design a custom kernel for W4A4 GEMV. We also integrate the de-quantization process into the GEMM and GEMV kernels. Table 9(b) presents the speedup ratios of the QuaRot kernel and our kernel compared to FP16. With a sequence length of 1, our quantized linear layer improves the QuaRot speedup from to on the RTX 3090, and from to on the A100-80GB. With a sequence length of 2048, our layer enhances the QuaRot speedup from to on the RTX 3090, and from to on the A100-80GB.
Appendix E Quantization Time
Table 10 shows the quantization time for PrefixQuant. PrefixQuant identifies prefixed tokens quickly, taking only 0.2 minutes for Llama-3-8B and 1 minute for Llama-3-70B. In contrast, the recent CushionCache (Son et al., 2024) requires 12 hours for the same task on Llama-3-8B. Additionally, the grid-search initialization is efficient, taking 0.7 minutes for Llama-3-8B and 12 minutes for Llama-3-70B. Experiments in Tables 3 and 4 demonstrate that PrefixQuant, even without fine-tuning, outperforms previous methods (Lin et al., 2024b; Ashkboos et al., 2024b). Fine-tuning requires more time, taking 2.2 hours for Llama-3-8B and 17 hours for Llama-3-70B, but it can successfully enhances the potential of low-bit quantization.
@@ Appendix E Quantization Time
Appendix F More Ablation Results
@@
| Fine-Tuning | Quant Type | W4A8KV4 | W4A4KV4 |
| No | token-wise dynamic | 6.84 | 8.29 |
| No | tensor-wise static | 6.90 | 7.93 |
| Yes | token-wise dynamic | 6.60 | 7.88 |
| Yes | tensor-wise static | 6.63 | 7.41 |
Dynamic Quantization. Tables 3 and 4 show that PrefixQuant with static quantization can surpass previous state-of-the-art methods (Xiao et al., 2023a; Ashkboos et al., 2024b; Lin et al., 2024b) with dynamic quantization. Note that without prefixed outliers, per-token dynamic quantization consistently outperforms per-tensor static quantization across different precisions, as shown in Table 6. Therefore, a question arises: can dynamic quantization further improve the performance of PrefixQuant? We replace per-tensor static activation quantization in PrefixQuant with per-token dynamic quantization and report the results in Table 13. We find that the winner varies with different precision. Specifically, per-token dynamic quantization marginally surpasses per-tensor static quantization in W4A8KV4 quantization, while per-tensor static quantization significantly outperforms per-token dynamic quantization in W4A4KV4 quantization. This is because, in high-precision quantization such as 8-bit, clipping is not necessary (Gong et al., 2024), and the MAX-MIN initialization of dynamic quantization adapts to a more diverse range flexibly. However, in low-precision quantization such as 4-bit, clipping is crucial to balance clipping error and rounding error (Lin et al., 2023), resulting in per-tensor static quantization outperforming per-token dynamic quantization.
@@
Appendix G Extend to Weight-only Quantization
@@
Appendix H Full Results of Weight-Activation quantization
@@
| PrefixQuant | per-tensor static | 5.48 | |
@@
| PrefixQuant | per-tensor static | 4.89 | |
@@
| PrefixQuant | per-tensor static | 3.39 | |
@@
| PrefixQuant | per-tensor static | 6.17 |
H.1 Comparisons with Related Works
@@
H.2 Detailed Accuracy Results
@@
H.3 Results on More Models
@@
| PrefixQuant w/o FT | W4A4KV4 | 67.80 | 73.75 | 43.94 | 71.51 | 77.2 | 66.84 | |
| PrefixQuant | W4A4KV4 | 66.54 | 73.42 | 43.09 | 71.17 | 77.64 | 66.37 | |
@@
| QuaRot | W4A8KV8 | 66.77 | 74.56 | 43.86 | 72.39 | 77.97 | 67.11 | |
| PrefixQuant w/o FT | W4A8KV8 | 69.14 | 75.12 | 44.45 | 73.06 | 77.53 | 67.86 | |
| PrefixQuant | W4A8KV8 | 69.06 | 75.25 | 44.8 | 73.19 | 78.13 | 68.09 | |
[…]
| QuaRot | W8A8KV8 | 68.98 | 75.96 | 46.59 | 74.41 | 79.11 | 69.01 | |
| PrefixQuant w/o FT | W8A8KV8 | 70.48 | 76.62 | 45.65 | 73.91 | 78.18 | 68.97 | |
@@
| PrefixQuant w/o FT | W4A4KV4 | 72.06 | 76.54 | 46.67 | 75.8 | 78.51 | 69.92 | |
| PrefixQuant | W4A4KV4 | 72.53 | 76.12 | 47.70 | 76.09 | 79.38 | 70.36 | |
@@
| QuaRot | W4A8KV8 | 70.24 | 78.21 | 47.01 | 74.49 | 79.87 | 69.96 | |
| PrefixQuant w/o FT | W4A8KV8 | 72.77 | 77.49 | 48.12 | 77.06 | 79.92 | 71.07 | |
| PrefixQuant | W4A8KV8 | 72.77 | 77.54 | 48.72 | 76.81 | 80.41 | 71.25 | |
[…]
| QuaRot | W8A8KV8 | 71.98 | 79.35 | 49.23 | 77.4 | 80.47 | 71.69 | |
| PrefixQuant w/o FT | W8A8KV8 | 72.53 | 78.38 | 48.98 | 76.81 | 80.9 | 71.52 | |
@@
| PrefixQuant w/o FT | W4A4KV4 | 75.45 | 80.51 | 52.3 | 77.06 | 81.12 | 73.29 | |
| PrefixQuant | W4A4KV4 | 77.35 | 82.3 | 56.4 | 79.29 | 82.05 | 75.48 | |
@@
| QuaRot | W4A8KV8 | 77.03 | 83.30 | 57.08 | 81.27 | 82.86 | 76.31 | |
| PrefixQuant w/o FT | W4A8KV8 | 77.35 | 82.79 | 54.35 | 78.28 | 82.21 | 75.00 | |
| PrefixQuant | W4A8KV8 | 79.08 | 83.56 | 57.42 | 80.39 | 82.05 | 76.50 | |
[…]
| QuaRot | W8A8KV8 | 77.82 | 83.8 | 57.34 | 80.93 | 82.75 | 76.53 | |
| PrefixQuant w/o FT | W8A8KV8 | 79.16 | 84.14 | 55.8 | 78.87 | 82.59 | 76.11 | |
@@
| PrefixQuant w/o FT | W4A4KV4 | 69.14 | 75.46 | 47.1 | 72.94 | 77.2 | 68.37 | |
| PrefixQuant | W4A4KV4 | 71.9 | 75.44 | 50.68 | 78.32 | 79.05 | 71.08 | |
@@
| QuaRot | W4A8KV8 | 72.74 | 77.35 | 51.62 | 77.48 | 79.22 | 71.68 | |
| PrefixQuant w/o FT | W4A8KV8 | 71.19 | 77.65 | 48.98 | 73.99 | 79.65 | 70.29 | |
| PrefixQuant | W4A8KV8 | 72.53 | 77.97 | 52.65 | 79.25 | 79.92 | 72.46 | |
[…]
| QuaRot | W8A8KV8 | 72.53 | 78.99 | 53.67 | 78.03 | 80.63 | 72.77 | |
| PrefixQuant w/o FT | W8A8KV8 | 74.11 | 79.25 | 53.75 | 78.03 | 80.36 | 73.10 | |
@@
| PrefixQuant w/o FT | W4A4KV4 | 77.43 | 83.48 | 58.87 | 79.88 | 82.32 | 76.40 | |
| PrefixQuant | W4A4KV4 | 77.35 | 83.79 | 60.15 | 81.31 | 83.3 | 77.18 | |
@@
| QuaRot | W4A8KV8 | 80.35 | 84.03 | 62.12 | 84.64 | 83.46 | 78.92 | |
| PrefixQuant w/o FT | W4A8KV8 | 79.23 | 84.71 | 59.39 | 81.57 | 84.22 | 77.82 | |
| PrefixQuant | W4A8KV8 | 79.48 | 84.86 | 62.29 | 82.53 | 84.33 | 78.70 | |
[…]
| QuaRot | W8A8KV8 | 80.66 | 84.84 | 63.65 | 85.56 | 84.44 | 79.83 | |
| PrefixQuant w/o FT | W8A8KV8 | 79.40 | 85.5 | 61.43 | 82.49 | 84.22 | 78.61 |
@@
| W4A8KV4 | 5.96 | 77.74 | 81.97 | 65.87 | 84.93 | 81.56 | 78.41 | |
| W4A4KV4 | 6.80 | 75.93 | 80.64 | 64.76 | 83.88 | 81.23 | 77.29 |
Appendix I More Visualizations
I.1 Outlier Token

### v2

@@ Appendix B Details of Rotation
Appendix C Quantization Time
Table 6 shows the quantization time for PrefixQuant. PrefixQuant identifies prefixed tokens quickly, taking only 0.2 minutes for Llama-3-8B and 1 minute for Llama-3-70B. In contrast, the recent CushionCache (Son et al., 2024) requires 12 hours for the same task on Llama-3-8B. Additionally, the grid-search initialization is efficient, taking 0.7 minutes for Llama-3-8B and 12 minutes for Llama-3-70B. Experiments in Tables 2 and 3 demonstrate that PrefixQuant, even without fine-tuning, outperforms previous methods (Lin et al., 2024b; Ashkboos et al., 2024b). Fine-tuning requires more time, taking 2.2 hours for Llama-3-8B and 17 hours for Llama-3-70B, but it can successfully enhances the potential of low-bit quantization.
@@ Appendix C Quantization Time
Appendix D More Ablation Results
@@
| Method | Precision | PPL. |
| - | FP16 | 5.54 |
| QuaRot | W4A8KV4 | 6.79 |
| DuQuant | W4A8KV4 | 6.19 |
| PrefixQuant-O1 | W4A8KV4 | 5.94 |
| PrefixQuant-O2 | W4A8KV4 | 6.04 |
| QuaRot | W4A4KV4 | 8.41 |
| DuQuant | W4A4KV4 | 7.27 |
| PrefixQuant-O1 | W4A4KV4 | 6.58 |
| PrefixQuant-O2 | W4A4KV4 | 6.82 |
Appendix E Comparisons in Long-Context Scenarios
PrefixQuant-O1 uses a shared clipping factor for each layer, while PrefixQuant-O2 further shares a scaling factor across the entire tensor. A longer input context implies that more activations share the same clipping or scaling factor. This raises a concern about whether PrefixQuant remains effective in long-context scenarios. Table 11 shows that PrefixQuant consistently outperforms existing methods at a context length of 8192, demonstrating the strong generalization ability of the proposed method.
@@ Appendix E Comparisons in Long-Context Scenarios
Appendix F Extend to Weight-only Quantization
@@
Appendix G Full Results of Weight-Activation quantization
@@
| PrefixQuant-O2 | per-tensor static | 5.48 | |
@@
| PrefixQuant-O2 | per-tensor static | 4.89 | |
@@
| PrefixQuant-O2 | per-tensor static | 3.39 | |
@@
| PrefixQuant-O2 | per-tensor static | 6.17 |
G.1 Comparisons with Related Works
@@
G.2 Detailed Accuracy Results
@@
G.3 Results on More Models
@@
| PrefixQuant-O1 w/o FT | W4A4KV4 | 66.85 | 74.27 | 43.86 | 72.35 | 77.97 | 67.06 | |
| PrefixQuant-O1 | W4A4KV4 | 67.48 | 73.77 | 43.17 | 71.3 | 77.97 | 66.74 | |
| PrefixQuant-O2 w/o FT | W4A4KV4 | 67.80 | 73.75 | 43.94 | 71.51 | 77.2 | 66.84 | |
| PrefixQuant-O2 | W4A4KV4 | 66.54 | 73.42 | 43.09 | 71.17 | 77.64 | 66.37 | |
@@
| QuaRot | W4A8KV4 | 66.77 | 74.56 | 43.86 | 72.39 | 77.97 | 67.11 | |
| PrefixQuant-O1 w/o FT | W4A8KV8 | 69.53 | 75.49 | 44.8 | 73.32 | 77.64 | 68.16 | |
| PrefixQuant-O1 | W4A8KV8 | 69.69 | 75.3 | 44.28 | 73.19 | 77.75 | 68.04 | |
| PrefixQuant-O2 w/o FT | W4A8KV4 | 69.14 | 75.12 | 44.45 | 73.06 | 77.53 | 67.86 | |
| PrefixQuant-O2 | W4A8KV4 | 69.06 | 75.25 | 44.8 | 73.19 | 78.13 | 68.09 | |
@@
| PrefixQuant-O1 w/o FT | W4A4KV4 | 72.38 | 76.92 | 47.7 | 75.8 | 79.43 | 70.45 | |
| PrefixQuant-O1 | W4A4KV4 | 71.98 | 76.3 | 46.59 | 76.43 | 78.94 | 70.05 | |
| PrefixQuant-O2 w/o FT | W4A4KV4 | 72.06 | 76.54 | 46.67 | 75.8 | 78.51 | 69.92 | |
| PrefixQuant-O2 | W4A4KV4 | 72.53 | 76.12 | 47.70 | 76.09 | 79.38 | 70.36 | |
@@
| QuaRot | W4A8KV4 | 70.24 | 78.21 | 47.01 | 74.49 | 79.87 | 69.96 | |
| PrefixQuant-O1 w/o FT | W4A8KV8 | 72.77 | 77.6 | 48.46 | 77.36 | 80.36 | 71.31 | |
| PrefixQuant-O1 | W4A8KV8 | 72.14 | 77.71 | 48.98 | 76.89 | 80.52 | 71.25 | |
| PrefixQuant-O2 w/o FT | W4A8KV4 | 72.77 | 77.49 | 48.12 | 77.06 | 79.92 | 71.07 | |
| PrefixQuant-O2 | W4A8KV4 | 72.77 | 77.54 | 48.72 | 76.81 | 80.41 | 71.25 | |
@@
| PrefixQuant-O1 w/o FT | W4A4KV4 | 77.98 | 81.38 | 55.55 | 78.7 | 81.12 | 74.95 | |
| PrefixQuant-O1 | W4A4KV4 | 78.77 | 83.23 | 56.48 | 79.92 | 82.75 | 76.23 | |
| PrefixQuant-O2 w/o FT | W4A4KV4 | 75.45 | 80.51 | 52.3 | 77.06 | 81.12 | 73.29 | |
| PrefixQuant-O2 | W4A4KV4 | 77.35 | 82.3 | 56.4 | 79.29 | 82.05 | 75.48 | |
@@
| QuaRot | W4A8KV4 | 77.03 | 83.30 | 57.08 | 81.27 | 82.86 | 76.31 | |
| PrefixQuant-O1 w/o FT | W4A8KV8 | 78.06 | 83.64 | 55.12 | 79.71 | 82.15 | 75.74 | |
| PrefixQuant-O1 | W4A8KV8 | 79.64 | 83.97 | 57.68 | 80.18 | 82.64 | 76.82 | |
| PrefixQuant-O2 w/o FT | W4A8KV4 | 77.35 | 82.79 | 54.35 | 78.28 | 82.21 | 75.00 | |
| PrefixQuant-O2 | W4A8KV4 | 79.08 | 83.56 | 57.42 | 80.39 | 82.05 | 76.50 | |
@@
| PrefixQuant-O1 w/o FT | W4A4KV4 | 70.32 | 75.86 | 48.38 | 71.46 | 77.58 | 68.72 | |
| PrefixQuant-O1 | W4A4KV4 | 70.88 | 75.95 | 52.47 | 78.7 | 79.6 | 71.52 | |
| PrefixQuant-O2 w/o FT | W4A4KV4 | 69.14 | 75.46 | 47.1 | 72.94 | 77.2 | 68.37 | |
| PrefixQuant-O2 | W4A4KV4 | 71.9 | 75.44 | 50.68 | 78.32 | 79.05 | 71.08 | |
@@
| QuaRot | W4A8KV4 | 72.74 | 77.35 | 51.62 | 77.48 | 79.22 | 71.68 | |
| PrefixQuant-O1 w/o FT | W4A8KV8 | 71.74 | 77.99 | 50.17 | 74.07 | 79.22 | 70.64 | |
| PrefixQuant-O1 | W4A8KV8 | 73.16 | 77.95 | 52.56 | 79.17 | 80.03 | 72.57 | |
| PrefixQuant-O2 w/o FT | W4A8KV4 | 71.19 | 77.65 | 48.98 | 73.99 | 79.65 | 70.29 | |
| PrefixQuant-O2 | W4A8KV4 | 72.53 | 77.97 | 52.65 | 79.25 | 79.92 | 72.46 | |
@@
| PrefixQuant-O1 w/o FT | W4A4KV4 | 77.74 | 83.61 | 58.19 | 80.3 | 82.43 | 76.45 | |
| PrefixQuant-O1 | W4A4KV4 | 77.74 | 84.06 | 58.96 | 81.31 | 83.35 | 77.08 | |
| PrefixQuant-O2 w/o FT | W4A4KV4 | 77.43 | 83.48 | 58.87 | 79.88 | 82.32 | 76.40 | |
| PrefixQuant-O2 | W4A4KV4 | 77.35 | 83.79 | 60.15 | 81.31 | 83.3 | 77.18 | |
@@
| QuaRot | W4A8KV4 | 80.35 | 84.03 | 62.12 | 84.64 | 83.46 | 78.92 | |
| PrefixQuant-O1 w/o FT | W4A8KV8 | 78.14 | 84.92 | 59.73 | 81.06 | 83.79 | 77.53 | |
| PrefixQuant-O1 | W4A8KV8 | 79.4 | 85.03 | 61.69 | 81.9 | 84.49 | 78.50 | |
| PrefixQuant-O2 w/o FT | W4A8KV4 | 79.23 | 84.71 | 59.39 | 81.57 | 84.22 | 77.82 | |
| PrefixQuant-O2 | W4A8KV4 | 79.48 | 84.86 | 62.29 | 82.53 | 84.33 | 78.70 |
@@
| W4A8KV4-O2 | 5.96 | 77.74 | 81.97 | 65.87 | 84.93 | 81.56 | 78.41 | |
| W4A4KV4-O1 | 6.49 | 75.77 | 81.51 | 65.96 | 83.8 | 81.28 | 77.66 | |
| W4A4KV4-O2 | 6.80 | 75.93 | 80.64 | 64.76 | 83.88 | 81.23 | 77.29 |
Appendix H More Visualizations
H.1 Outlier Token